RYDBERG STATE SPECTROSCOPY OF BORON HYDRIDE RADICALS.

\underline{JASON CLARK} and EDWARD R. GRANT, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

Ionization detected electronic absorption spectra of radical boron hydrides will be reported. BH_{x} radicals are produced by ArF (193 nm) photolysis of B_{2}H_{6} in a supersonic pulsed-jet. Ions are mass-resolved and detected using a time-of-flight mass spectrometer. Various Rydberg states of BH and BH_{2} will be discussed.